OpenStudy (anonymous):

laplace transform: L{f(t)}? = -1, 0<=t<1 0, 1<=t<3 1, t>=3

OpenStudy (anonymous):

Piecewise Function?

OpenStudy (anonymous):

\[\int _0^1-1 e^{-s t}dt\text{ }+ \int _1^30 * e^{-s t}dt +\int _3^{\infty }1 e^{-s t}dt\]

OpenStudy (anonymous):

I got \[\frac{e^{-3 s}}{s}+\frac{-1+e^{-s}}{s}\] not sure if that's right though

OpenStudy (anonymous):

your answer spark something..haha..i'm trying too..hopefully its the same!..tx..:)

OpenStudy (anonymous):

yes, that's what I got using computer too

OpenStudy (anonymous):

i couldnt get the -1/s

OpenStudy (anonymous):

haha...got it!..tx a lot!..:)

OpenStudy (anonymous):

great, btw what class is it for?

OpenStudy (anonymous):

differential equation. :)

OpenStudy (anonymous):

me too, have you done fourier series yet?

OpenStudy (anonymous):

done. its the final chpter in this subjct, btw im having final paper on ODE this coming wednesday..:X

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

tx again.